40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1561     Entry time: Fri May 8 02:39:02 2009
Author: pete, rana 
Type: Update 
Category: Locking 
Subject: crossover 

attached plot shows MC_IN1/MC_IN2.  needs work.

This is supposed to be a measurement of the relative gain of the MCL and AO paths in the CM servo. We expect there to

be a more steep slope (ideally 1/f). Somehow the magnitude is very shallow and so the crossover is not stable. Possible

causes? Saturations in the measurement, broken whitening filters, extremely bad delay in the digital system? needs work.

 

Attachment 1: crossover.pdf  5 kB  | Hide | Hide all
crossover.pdf
Attachment 2: photo.jpg  139 kB  Uploaded Fri May 8 04:58:15 2009  | Hide | Hide all
photo.jpg
ELOG V3.1.3-